OpenStudy (anonymous):

Expand logarithm: e^(5lnx)+ln1

OpenStudy (zzr0ck3r):

\(e^{5\ln(x)}+\ln(1)=e^{\ln(x^5)}+0=x^5 \ \\\forall x>0\)

OpenStudy (zzr0ck3r):

that second line just says for all x>0 note that we cant put \(x\le 0\) because we can only put positive numbers into \(\ln(x)\) so they are only truly equal if we make that note